### v1

# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust

Abstract.
In recent years, various methods have been introduced for explaining the outputs of “black-box” AI models. However, it is not well understood whether users actually comprehend and trust these explanations. In this paper, we focus on explanations for a regression tool for assessing cancer risk and examine the effect of the explanations’ content and format on the user-centric metrics of comprehension and trust. Regarding content, we experiment with two explanation methods: the popular SHAP, based on game-theoretic notions and thus potentially complex for everyday users to comprehend, and occlusion-1, based on feature occlusion which may be more comprehensible. Regarding format, we present SHAP explanations as charts (SC), as is conventional, and occlusion-1 explanations as charts (OC) as well as text (OT), to which their simpler nature also lends itself. The experiments amount to user studies questioning participants, with two different levels of expertise (the general population and those with some medical training), on their subjective and objective comprehension of and trust in explanations for the outputs of the regression tool. In both studies we found a clear preference in terms of subjective comprehension and trust for occlusion-1 over SHAP explanations in general, when comparing based on content. However, direct comparisons of explanations when controlling for format only revealed evidence for OT over SC explanations in most cases, suggesting that the dominance of occlusion-1 over SHAP explanations may be driven by a preference for text over charts as explanations. Finally, we found no evidence of a difference between the explanation types in terms of objective comprehension. Thus overall, the choice of the content and format of explanations needs careful attention, since in some contexts an explanation’s format, rather than its content, may play the critical role in improving user experience.
Explainable AI, Feature Attribution, User Studies, Trust
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Explainable AI (XAI) has risen to prominence in recent years as a means of giving users the explanations they need to be able to trust the predictions of “black-box” AI models. This trustworthiness of AI models is especially pertinent in high risk settings such as medicine, when patients’ lives may depend on these predictions. The ways in which explainability may facilitate trust in users has been studied extensively (Jacovi et al., 2021; Schoeffer et al., 2022; Ferrario and Loi, 2022; Panigutti et al., 2022), with the general consensus leaning towards it playing an important role in this critical bond between humans and machines. However, despite this growing consensus, what remains to be seen is which forms of XAI, and which requisite characteristics, are best suited to facilitating trust in users in each individual application of AI models. One of the most prominent types of XAI method is feature attribution (FA, see (Bodria et al., 2023) for an overview). FA explanations assign each feature in a given input a positive or negative attribution value denoting its importance towards the AI model’s output. However, blindly applying one FA method or another in a particular situation is not a scientifically rigorous approach to accounting for the human factors involved in enabling trustworthy AI.
The evaluation of XAI methods, particularly those concerning (the often neglected (Keane et al., 2021)) user studies with humans, have thus been of critical importance to ensuring the methods are serving humans’ needs. Various metrics have been studied as dimensions along which to make these evaluations, including those which are machine-centric, e.g. robustness (Jiang et al., 2023), i.e. how sensitive an explanation is to changes in the scenario. However, it is human-centric metrics which present direct opportunities for measuring how humans feel about XAI, such as (perceived) comprehensibility (Schmude et al., 2023), how well users (think they) understand explanations, and trust (Jacovi et al., 2021), i.e. how much users trust explanations. We argue that each individual application setting requires its own analysis along these metrics and optimisation with regards to its unique explanatory requirements, as is called for in (Hamon et al., 2021). Further, very few studies, with some notable exceptions, e.g. (Szymanski et al., 2021; Bertrand et al., 2023; Schmude et al., 2023), have considered the effect of the format, rather than the content, of explanations on these metrics.
Concurrently, the application of AI models to the medical domain has been swift and seemingly ubiquitous (see (Tjoa and Guan, 2021) for a recent overview), with suitable areas ranging from image classification for brain tumour detection (Legastelois et al., 2023) to clinical decision support (Domínguez et al., [n.d.]). The huge potential for tangible benefits of AI models to countless human lives means the search for techniques to ensure that the AI models are trustworthy is particularly urgent, and thus XAI methods are in increasing demand here. One recent study (Kostopoulou et al., 2022) investigates whether predictions from a cancer risk prediction model, namely the QCancer algorithm (Hippisley-Cox and Coupland, 2015), have a beneficial effect in supporting the referral decisions of UK-based General Practitioners (GPs). Kostopoulou et al. (2022) found that the inclination of the GPs on whether to refer the patient changed 26% of the time, while their decisions switched entirely only 3% of the time. However, despite the reluctance of GPs to adopt advice from the algorithm, the changed decisions were more consistent with established clinical guidelines, highlighting the potential of this approach to decision support. In a follow-up study, Pálfi et al. (2023) attempted to improve the uptake of the algorithm by providing a naive, chart-based FA explanation that visualised the relative weights (based on the regression coefficients of the model) of the features for each prediction. However, they found no additional benefit of the explanations over providing the prediction without the explanation. They speculated that this failure to enhance the adoption of and trust in the algorithm by providing an XAI method could be related to the difficulty in understanding the explanations. This context thus seems to be relatively fertile ground to examine the potential of established XAI techniques. However, given that the explanatory requirements of AI models within healthcare have been shown to be different from those in general domains (Legastelois et al., 2023), user studies which are tailored to individual contexts are needed, as stated in (Colin et al., 2022).
In this paper, we set out to address some of the unanswered questions outlined above, in particular those highlighted by Pálfi et al. (2023). Namely, we aimed to shed light on differing degrees of (perceived) comprehensibility and trust affected in users when they receive explanations of different types and formats. We do this by deploying different types of XAI methods, namely the FA explanations SHAP (Lundberg and Lee, 2017) and occlusion-1 (Ancona et al., 2018), for the QCancer algorithm, in the realm of prediction algorithms for diagnosis in healthcare. We then undertook two user studies, each with a different group of participants, one sampled from the general population and one from students in medical training, to examine the effect of different users’ explanation goals (Delaney et al., 2023), which has been shown to be an important factor in the evaluation of explanations (Kim et al., 2023). The participants were asked to answer questions in a survey to assess their perceived comprehensibility of and trust in explanations for the outputs of the QCancer algorithm on fictitious patients described by vignettes from (Pálfi et al., 2023). Moreover, we created a novel objective measure of comprehension to control for any potential reporting biases. Differently from existing studies, we also control for the format of explanations (chart or text) to examine their effect on this paradigm. It should be noted, however, that we do not examine text-based SHAP explanations since describing their game-theoretic notions in full would be incomprehensible to most users, and any approximation would be unrepresentative and unfair (we expand on the reasoning for this decision in §3.2). As such, our chosen research questions (RQs) are as follows, all considered within this medical context:
RQ1: Are occlusion-1 explanations, in a.) chart or b.) text format, more comprehensible than SHAP explanations?
RQ2: Are occlusion-1 explanations, in a.) chart or b.) text format, perceived as being more comprehensible than SHAP explanations?
RQ3: Are occlusion-1 explanations, in a.) chart or b.) text format, perceived as being more trustworthy than SHAP explanations?
Our expectations were that the null hypothesis would be rejected in all cases (i.e. the RQs would be answered affirmatively), given the additional complexity in SHAP explanations. We were surprised to discover instead that the dominance of occlusion-1 over SHAP explanations along these metrics is unclear and, where present, was possibly due to the use of text format.
The paper is structured as follows. We first overview the literature which is most relevant to our work (§2), then detail the methodology of our studies (§3) and our findings therefrom (§4). We then discuss the results and state the limitations of the studies, before concluding and looking ahead to future work5).
2. Related Work
Of the multitude of works undertaking user studies comparing different types of XAI, one which most closely aligns with ours is that of Bertrand et al. (2023). They deliver FA explanations in different formats to examine textual explanations vs. graphical explanations. They find that the former increases trust relative to the latter, but they posit that this could be due to the explanations’ simplicity. Somewhat in contradiction with other work (Rago et al., 2021), they found that explanations did not improve users’ understanding of models’ recommendations.
Other works made similar conclusions regarding the format of explanations, e.g. Szymanski et al. (2021) assess visual and textual explanations for the predictions of AI models in user comprehensibility, controlling for the expertise of users, as we do in our studies. The authors found that lay users preferred the visual explanations but could easily misinterpret them compared to textual explanations.
Another closely related study (Schmude et al., 2023) evaluates the effect of textual, dialogical and interactive explanations on user understanding. They found that the textual explanations presented greater limits to understanding for and lower preferences from users.
Moving away from the format of explanations, Cheng et al. (2019) compared white- vs. black-box explanations. The authors assessed users’ (perceived) understanding and trust, as we do, validating our decision to differentiate between self-reported and objective comprehensibility. Cheng et al. (2019) found that, among other results related to interactivity, white-box explanations did not increase understanding relative to those which are black-box, validating our decision to use the latter. Surprisingly, however, they also found that education or technical literacy had no effect on these parameters. However, Schoeffer et al. (2022) found that more information gives greater perceptions of information fairness and trustworthiness and that AI literacy also increased both perceptions.
The transparency of explanations is also studied by Bove et al. (2022), who investigate the effect on AI models’ explanations of contextualisation, i.e. giving more transparency by showing what the system does and how it does it, and exploration, i.e. interacting with the explanation/model or navigating different explanations. They assess end users’ understanding of and satisfaction with explanations, failing and succeeding, respectively, to reject the null hypotheses.
Hadash et al. (2022) identify a relevant issue with FA explanations concerning the negative framing of attributions, i.e. the case where a feature with a negative attribution contributes towards a positive decision or vice versa. A user study shows that the positive framing of feature attributions gives significant benefits in terms of user comprehensibility. In this paper we keep the FA explanations in their original framing but plan to investigate positive framing in the future.
Wang and Yin (2023) find that changes to explanations (i.e. lacking robustness) may reduce users’ trust in and satisfaction with the AI model but exclusively in the case where participants have domain knowledge. This demonstrates the importance of separating users by their expertise, as we have done in this paper. Users’ contexts, in the form of their personal relevance, AI expertise and level of stake, were also found to be important factors in designing XAI to inspire trust in the user study of (Yuan et al., 2023). This is backed up by other claims in the literature, e.g. that explanations must be designed to suit the needs of individuals (Bhatt et al., 2020) and also make sense to them (Byrne, 2023).
In summary, there are a number of orthogonal studies to those we undertake here, with many design choices validating those we take, but none consider the assessment of (perceived) comprehensibility and trust of FA explanations, controlling for the explanation format and the participant expertise.
Explainability is often considered within the medical domain as a potential tool to enhance trust in and uptake of prediction models (Tjoa and Guan, 2021; Watson et al., 2019). For instance, Lebovitz et al. (2022) conducted a user study in a hospital where AI tools were deployed by radiologists making predictions concerning breast cancer, lung cancer and bone age: they found that trust in the AI models’ predictions, and thus incorporation into the radiologists’ judgements, were highly dependent on their department, pointing to XAI as a potential solution.
In a user study with healthcare providers, Panigutti et al. (2022) assessed the effect of explanations on trust in AI models operating a Clinical Decision Support System. They showed that users preferred explanations from the AI model, though the user satisfaction with the explanations themselves was low. In contrast, a vignette study with intensive care doctors found no additional benefit of an FA explanation over a prediction only (Nagendran et al., 2023). Moreover, doctors, on average, reported somewhat negative attitudes about the usefulness of the explanations presented during the experiment (mean of 1.97 for non-training doctors on a scale from 1 to 5 with 5 being very useful).
In summary, there are few studies that test the potential of XAI methods to improve medical decisions, and they offer mixed results. Several recent reviews (Antoniadi et al., 2021; Nor et al., 2021; Hauser et al., 2022) highlighted these issues and called for further rigorous, empirical evaluation of the potential of XAI methods in healthcare, giving further motivation for our study.
3. Methodology
In this section, we detail our methodology for the user studies. We first cover the collection of the vignettes presented to participants with the predictions made by the QCancer algorithm (§3.1), before defining explanations for the predictions (§3.2). We then detail the experimental design and procedure (§3.3), explain how we recruited participants from the two different populations (§3.4) and outline our method for analysing the data (§3.5).
3.1. Vignettes
We adapted twenty fictitious vignettes (including one practice vignette) from (Pálfi et al., 2023), describing hypothetical patients presenting at the primary care (i.e. their GP) with symptoms and risk factors suggestive of gastro-oesophageal cancer. Note that we chose to focus on one type of cancer here for simplicity, though the QCancer algorithm and the original vignettes concern multiple types of cancer. Each vignette begins with a table including demographics and some risk factors of the patient, and continues with a textual description of a patient‘s symptoms, other risk factors and the results of some relevant examinations. All of the presented factors were predictive of gastro-oesophageal cancer according to the QCancer algorithm. Then, we provided the estimate of the patient’s risk of gastro-oesophageal cancer calculated by the QCancer algorithm followed by an explanation of this prediction and information on how the explanation should be interpreted. An example of one such vignette, with the output of the QCancer algorithm shown for gastro-oesophageal cancer is given in the following (we will give examples of the explanations in §3.2).
Example 0.
Patient name: Joyce Acorn (female), Age: 65, BMI: 30.64 (Height: 177cm, Weight: 96kg), Smoking: Smoker 20 cigarettes per day, Past medical history: Nil, Family history: Nil.
Joyce comes to see you about the indigestion she’s been experiencing over the past month. She also mentions that it has become increasingly difficult to swallow during meals. She enjoys food usually but has noticed that she is now off her food. She denies any post-menopausal bleeding. When she saw the practice nurse last week for her flu vaccine, she had also taken some routine blood tests. There are no other symptoms and examination findings are normal.
The cancer prediction algorithm estimates the risk of gastro-oesophageal cancer to be 31.90%.
The variables used in the QCancer algorithm’s prediction of gastro-oesophageal cancer and the values which Example 3.1 results in are shown in Table 1 (top and bottom rows, respectively – the table also shows, in the middle, the baseline patient values for males and females, which will be discussed in §3.2).
Gender Age BMI SS T2 CP LA WL DS AP BV I H A BH
Baseline Male M 40 25 non no no no no no no no no no no no
Baseline Female F 40 25 non no no no no no no no no no no no
Example 3.1 F 65 30.64 heavy no no yes no yes no no yes no no no
Table 1. Values for the QCancer algorithm used for the baseline patients and the patient in Example 3.1, with variables including Smoking Status (SS), Type 2 Diabetes (T2), Chronic Pancreatitis (CP), Loss of Appetite (LA), Unintentional Weight Loss (WL), Difficulty Swallowing (DS), Abdominal Pain (AP), Blood in Vomit (BV), Indigestion (I), Heartburn (H), Anaemia (A) and Change in Bowel Habit (BH).
The vignettes were divided into a low (less than 2%) and high (more than 2%) risk sets based on the estimated risk of gastro-oesophageal cancer for the patient computed by the QCancer algorithm. The low risk set ranged from 0.08% to 1.49%, and in the high risk set they ranged from 2.53% to 53.45%. To control for order effects and to avoid the repetition of vignettes during the experiment, the sets of low and high risk vignettes were further divided into two subsets of vignettes with similar risk profiles, giving low risk sets A and B, and high risk sets A and B (see more details in §3.3).
3.2. Explanations for the QCancer Algorithm
In this section, we define the explanations which we deliver to users. To define these explanations we assume a simple, single-label classification problem comprising a classifier $f:\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}\rightarrow[0,1]$, where $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$ represent the domains for each of the $k$ features ($k\!\!=\!\!15$ in our setting, see Table 1). Then, for any input $\mathbf{x}\!\!\in\!\!\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{k}$, $f(\mathbf{x})\!\!\in\!\![0,1]$ is the predicted probability of the class given the current values of the input features. This represents the QCancer algorithm, where the features, e.g. age, are inputs to the algorithm and the output is the predicted risk of gastro-oesophageal cancer. We refer to the $i$th value of any input $\mathbf{x}$ as $x_{i}$, where $i\!\!\in\!\!\{1,\ldots,k\}$. In order to define the FA explanations, we first introduce a simplified version of any given input $\mathbf{x}$ as $\mathbf{x}^{\prime}$, where there exists a mapping function $h_{\mathbf{x}}$ such that $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$, i.e. returning the original input given the simplified version. We then let, as in (Lundberg and Lee, 2017), $\mathbf{x}^{\prime}\!\!\in\!\!\{0,1\}^{M}$, where $M$ is the number of simplified input features (note that $M\!\!=\!\!k$ throughout this paper). Then, if $x_{i}^{\prime}\!\!=\!\!1$, then $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the current value of the feature, while if $x_{i}^{\prime}\!\!=\!\!0$, then $\mathbf{x}\!\!=\!\!h_{\mathbf{x}}(\mathbf{x}^{\prime})$ sets $x_{i}$ to the feature’s baseline value, i.e. a pre-defined default value. We set the baseline values (referred to as baseline patients in the explanations) to one of two sets of values representing healthy males and females depending on the patient’s gender, as indicated in Table 1, which were used for the studies.
One of the most prominent of all FA methods is SHAP (Lundberg and Lee, 2017), where each feature of an input to a model is assigned an importance value computed from a game-theoretical approximation of its contribution.
Definition 3.2.
Given a classifier $f$ and an input $\mathbf{x}$, the SHAP explanation for $x_{i}$ wrt $f$ is:
$e_{S}(f,\mathbf{x},i)=\sum_{\mathbf{z}^{\prime}\subseteq\mathbf{x}^{\prime}}{\frac{|\mathbf{z}^{\prime}|!(M-|\mathbf{z}^{\prime}|-1)!}{M!}f(h_{\mathbf{x}}(\mathbf{z}^{\prime}))-f(h_{\mathbf{x}}(\mathbf{z}_{-i}^{\prime}))}$
where $\mathbf{z}^{\prime}\subseteq\mathbf{x}^{\prime}$ represents that $\mathbf{z}^{\prime}$ is a vector whose non-zero entries are a subset of the non-zero entries in $\mathbf{x}^{\prime}$, $|\mathbf{z}^{\prime}|$ is the number of non-zero entries in $\mathbf{z}^{\prime}$ and $\mathbf{z}_{-i}^{\prime}$ denotes $\mathbf{z}^{\prime}$ after setting $z_{i}^{\prime}=0$.
For a given input, SHAP assigns each feature an attribution value based on a weighted sum considering the effect of changing the feature from its current value to the baseline value in a range of combinations of other features’ values.
We then introduce chart-based SHAP (SC) explanations, which are the explanations computed by SHAP expressed as a percentage in chart format, as exemplified in Figure 1 (bottom). [^1] On acceptance, we will share the link to our explanation visualisation tool. Note that, for all explanations, the displayed features were the six with the highest magnitudes. Note also that all explanations were accompanied by a textual description of the explanation, as shown in Figure 1 (top). In the example explanation in Figure 1, we can see that the top five features increased the risk of gastro-oesophageal cancer, with difficulty swallowing and its value of yes being the most important. The last feature, BMI, with value 30.643, reduced the risk but had a smaller magnitude, and thus importance, as calculated by SHAP.
Figure 1. A SC explanation (bottom) for the vignette given in Example 3.1, and textual description of the SC explanation (top). Red (blue) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
We then selected the occlusion-1 method (Ancona et al., 2018) for its simplicity in hopes of achieving more comprehensibility and trust in users. This method takes the set of baseline feature values (as in the SHAP explanations) and assigns any feature an attribution corresponding to the difference in the prediction if the feature’s value is changed to the baseline value (with all other features’ values remaining unchanged).
Definition 3.3.
Given a classifier $f$ and an input $\mathbf{x}$, the occlusion-1 explanation for $x_{i}$ wrt $f$ is:
$e_{O}(f,\mathbf{x},i)=f(h_{\mathbf{x}}(\mathbf{x}^{\prime}))-f(h_{\mathbf{x}}(\mathbf{x}_{-i}^{\prime}))$
where $\mathbf{x}_{-i}^{\prime}$ denotes $\mathbf{x}^{\prime}$ after setting $x_{i}^{\prime}=0$.
Note that, while these explanations do not explicitly mention datapoints, each feature’s attribution in explanations implicitly refers to a modified datapoint with that feature set to the baseline value, with all other features remaining unchanged. Thus, while this differs from counterfactual and semi-factual explanations in the literature (see (Aryal and Keane, 2023; Guidotti, 2022), respectively, for recent overviews [^2] Also that there are works which draw important links between FA and counterfactual explanations (Mothilal et al., 2021; Albini et al., 2022).), which are based on single modified datapoints, it still gives an actionable representation of (six) modifications users could make to the input to achieve a different output (similarly to Albini et al. (2020)). We therefore tentatively predicted that this inherent counterfactual nature of occlusion-1 explanations may make them more amenable to humans, as observed for counterfactual explanations by Miller (2019), which strengthened our prediction of the rejection of the null hypotheses for RQ1-3.
This definition thus allowed us to introduce two different formats of explanation. Firstly, chart-based occlusion-1 (OC) explanations are the explanations computed as in Definition 3.3 in chart format, as shown in Figure 2 with an accompanying textual description. We can see here that the features are the same as those selected by the SC explanations but the values are significantly different. However, the counterfactual reading of these explanations also allowed us to describe the attributions intuitively to users. Thus, we used the same occlusion-1 explanations, this time in textual format, to define text-based occlusion (OT) explanations, which are delivered as shown in Figure 3 with their accompanying textual description. This distinction between the two types of occlusion-1 explanation allows us to assess the effects not only of the content of an explanation, but also of its format. Note that SHAP allowed no such intuitive reading: the complexity of SHAP explanations in Definition 3.2 demonstrates they how do not lend themselves to text in what we might have called “ST” explanations, and that most users would not understand these game-theoretical notions. Indeed, it is difficult to picture, at least in a cognitively manageable form, what the ST equivalent for the SC explanation in Figure 1 would be, as the OT explanation in Figure 3 is to the OC explanation in Figure 2. We could have given an approximation of these explanations, but this simplification could be seen as being unrepresentative of SHAP, and it would be an unfair test. Thus, we refrained from defining and experimenting with this form of explanation.
Figure 2. An OC explanation (bottom) for the vignette given in Example 3.1, and textual description of the OC explanation (top). Red (green) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
Figure 3. An OT explanation (bottom) for the vignette given in Example 3.1, and textual description of the OT explanation (top). Note that the textual descriptions of the OT explanations were identical to those of the OC explanations.
3.3. Experimental Design & Procedure
We assessed the RQs in two different samples: a sample from the general population where a limited number of participants had medical experience (Study 1) and an expert sample where participants had some medical training and experience (Study 2). We created two surveys using Qualtrics [^3] https://www.qualtrics.com that followed the same structure, as follows.
  • Pre-Test (6 or 8 questions) – The participants were shown information about the study and were asked to provide consent. They were then asked questions regarding their background and experience of AI and cancer prediction algorithms (additional demographics questions were included for Study 2), including their trust in AI models in healthcare, before the task was explained to them with an illustrative example.
  • Main Test (4 x 6 questions) – In the main task, we provided the participants with four vignette sets, each containing a vignette, a prediction and an explanation (with its accompanying textual description), along with six questions on the explanation. The four sets concerned two types of explanation for comparison, each with a low risk and a high risk vignette. The six questions consisted of:
    • an attention check comprising the most important feature task (attention check; free text variable)
    • a perceived comprehensibility rating (addressing RQ2; interval scale from 0 to 10)
    • a trust rating (addressing RQ3; interval scale from 0 to 10)
    • a request for reasoning for the trust rating (optional; free text variable);
    • a request for additional feedback (optional; free text variable)
    • a comprehensibility test comprising the definition recognition task (addressing RQ1; categorical variable with four options).
  • Post-Test (4 questions) – The participants were asked to leave feedback if they had any technical difficulties during the survey, if they found the reimbursement fair and if they had any general comments about the study.
After each vignette, the participants were provided with an the attention check to assess whether they were engaging with the experiment. The attention check consisted in asking them which was the most important feature in determining the gastro-oesophageal cancer risk according to the explanation, with the correct answer being the feature which gave the highest absolute value of $e_{S}$ for SC explanations and $e_{O}$ for OC and OT explanations. For the perceived comprehensibility rating addressing RQ2 and the trust rating addressing RQ3, we asked the participants how well they thought they understood and how much they would trust, respectively, the explanation, on a scale of 0-10. Next, we asked for free-form textual feedback from the participants on reasons for their trust rating and on any additional points, both of which were optional. Finally, on a new page, we measured the participants’ comprehension of the explanation, addressing RQ1, by offering them four differently worded descriptions of the explanation as indicated below, one of which was correct for each explanation (as indicated in square brackets, but not shown to participants), and asked them which they believed it to be correct.
What would you say best describes a feature’s percentage indicated in the explanation on the previous page?
$\bullet$ The change in predicted risk when the feature’s current value is modified to the baseline value. [Correct for OC and OT explanations.]
$\bullet$ A representative importance of this feature’s value relative to the baseline value, based on different configurations of values for other features. [Correct for SC explanations.]
$\bullet$ The portion of the predicted risk associated with that feature.
$\bullet$ The average predicted risk for a patient with the feature’s current value.
Figure 4 illustrates the experimental design of Studies 1 and 2, which were identical in this regard. The participants were split such that 50% assessed SC vs. OC explanations (i.e. addressing RQ1a, RQ2a and RQ3a) and 50% assessed SC vs. OT explanations (i.e. addressing RQ1b, RQ2b and RQ3b). We then split the twenty vignettes into one vignette for demonstration to all participants, ten vignettes presenting low risk patients (less then 2% chance of gastro-oesophageal cancer) and nine vignettes presenting high risk patients (equal to or greater than 2% chance of gastro-oesophageal cancer), to ensure that all participants received one low and one high risk for each of the explanations that they examined. We then further split the high and low risk sets into two (giving High Risk Set A, High Risk Set B, Low Risk Set A and Low Risk Set B) so that the vignettes were presented to the participants evenly. Further, the four vignette sets for each participant were presented in random order to ensure there was no ordering bias.
Figure 4. Experimental design for user studies, with the two sets of high risk vignettes represented by shades of red, and the two sets of low risk vignettes represented by shades of green.
3.4. Participant Recruitment and Ethics Statement
For Study 1, we aimed to recruit 100 participants from the general population, who were over 18 and spoke English, via the online research platform Prolific [^4] Prolific (https://www.prolific.com) was chosen because of its superior data quality when compared with other platforms (Peer et al., 2022).. Upon completion, the participants were paid an average of \$4.50 for their time based on the platform’s standard reimbursement rates of \$21.20 per hour. Data collection was anonymous during Study 1. In Study 2, we aimed to recruit at least 50 students in medicine from UK-based universities. The potential participants were approached via student mailing lists, and were invited via a link to the survey in an email, which also contained the participant information sheet with details about the study. Note that all students were 18 years of age and English-speaking since they were attending university in the UK. At the end of Study 2, participants were asked to provide their university email address, to which we sent a £10 voucher for the online retailer Amazon to reimburse them for their time. All the participant responses were pseudononymised once this payment was made (i.e. we deleted their e-mail addresses from the dataset). Participation in both studies was voluntary, participants were asked to provide informed consent after reading the participant information sheet and they could withdraw at any time prior to completing the survey, simply by closing the questionnaire window, with their data deleted permanently. We informed the participants that once they had completed the survey, they could no longer withdraw and we would keep their data. [^5] The anonymised and pseudononymised data (respectively) will be shared via Zenodo (https://zenodo.org) on acceptance. We received ethics approval from our institution for these studies.
3.5. Data Analysis
Creation of the Main Variables of Interest
We created categorical variables with two levels from the responses to the most important feature and the definition recognition tasks, where each response was categorised as either correct or incorrect. When more than one feature was named as the most important, we categorised the response as correct as long as the first mentioned feature was the one with the strongest absolute contribution. Moreover, for the vignettes where the feature with the highest absolute contribution was a protective factor (i.e. the reduced risk of cancer), we accepted responses as correct where the strongest negative contributing feature was named first. We used the subjective understanding of the explanations and the trustworthiness of the explanations variables as interval variables.
Statistical Analyses
We ran the analyses in R (version 4.3.1). We built generalised linear and linear mixed-effects regression models with the lme4 R package (Bates et al., 2015) to analyse our categorical and interval variables, respectively. For the generalised linear regression models, we could only include random intercepts by participants (the models did not converge with more random intercepts), and, for the linear regression models, we included random intercepts by participants and vignettes. For the generalised linear and the linear regression models we reported Odds Ratios (ORs) and raw regression slopes (bs), respectively. We also reported 95% CIs for each regression coefficient. To test our hypotheses, we used the conventional p-value threshold of .05.
4. Results
Study Samples
Study 1 was completed by 99 participants [^6] One participant skipped the survey in Study 1 so we deleted their data and did not include them in the results., from which 1 reported to have colour-blindness (we did not measure demographics). Study 2 was completed by 69 medical students [^7] One participant left a non-university email in Study 2 so we deleted their data and did not include them in the results., from which 3 reported to have colour-blindness. The mean age of the participants in Study 2 was 22.3 (SD = 3.3), with 35 females (32 males, 1 other and 1 preferred not to disclose). As expected, participants in the medical student sample (Study 2) reported to have heard more about medical algorithms than participants in the Prolific sample (Study 1), but the number of people who had tried a medical AI tool before was negligible in both samples. Interestingly, trust towards algorithms in general was comparable in the two samples, and, on average, participants in both samples had neutral attitudes (roughly 5 on a scale from 1 to 9, where 5 was neutral). Table 2 presents the descriptive statistics broken down by samples. Since our sample sizes were defined by resource constraints, we conducted sensitivity power analyses as per the recommended good practice (Lakens, 2022). The multilevel regressions directly comparing the SC with either the OC or the OT explanations (RQ2 and RQ3) were powered at the 80% level to detect effect sizes of Cohen‘s d of 0.31 and 0.38, in Study 1 and 2, respectively. [^8] We used GPower 3.1 for the sensitivity power analyses, and we calculated the effective sample sizes for each study by multiplying the number of participants by the cluster sizes (i.e. each user responded to two vignettes) and then dividing it by the Design effect, which is 1+ (cluster size - 1) multiplied by the intra-class correlation (which we assumed to be low, r=0.2). Hence, we applied a Design effect of 1.2 for our studies.
Variable Response Study 1 (N = 99) Study 2 (N = 69)
Familiarity with Unfamiliar 85 (86%) 41 (59%)
AI (count) Not used 12 (12%) 26 (38%)
Used 2 (2%) 2 (3%)
Mean 4.84 5.20
Trust (0-10) Median 5 5
SD 2.17 2.00
Table 2. Descriptive statistics of the familiarity and trust in medical AI variables broken down by study samples.
Attention Check
We analysed the data from the most important feature task to assess whether the participants engaged with the experiment. Since all of our low risk vignettes used the same scaling for the chart-based explanations, our SC explanations were visualised with systematically smaller bars than the OC explanations, which may have created unintended difficulty in differentiating the bars on the charts for the former. Therefore, we calculated the percentages of correctly identified most important features separately for low and high risk cases, as well as for the explanations (See Table 3). Generally, the participants managed to correctly pinpoint the most important features in both studies, except for the chart-based explanations of the low risk cases in Study 1, where they failed in around half of the instances. These findings are in line with the notion of participants understanding and engaging with the task at hand but struggling to correctly respond when the task is more difficult.
Explanation High Risk Low Risk
Study 1 Study 2 Study 1 Study 2
SC 80.8% 92.8% 45.5% 69.6%
OC 85.1% 84.4% 57.4% 87.5%
OT 88.5% 94.6% 90.3% 89.2%
Table 3. Percentages of correctly identified most important features (Attention Check).
Objective Understanding of Explanations (RQ1)
We merged the OC and OT explanations to a single occlusion explanation group, since format was not relevant for the definition recognition task. First, we ran two intercept only regression models to compare performance in recognising the SHAP and occlusion definitions against chance level. In all cases, except for the occlusion explanations in Study 2, the 95% CIs of the ORs included chance level (OR of 1/3) suggesting that the participants struggled to recognise the correct definition of the explanations. Next, we analysed all data in a single regression model and entered explanation type as a predictor. We found no evidence that participants were better at recognising the definition of occlusion than SHAP explanations. Table 4 presents the corresponding ORs, their 95% CIs and the multilevel regressions.
Regression Model Study 1 Study 2
OR p OR p
SHAP 0.48 [0.28, 0.80] - 0.32 [0.12, 0.87] -
Occlusion 0.33 [0.16, 0.68] - 0.64 [0.38, 1.08] -
SHAP vs. Occlusion 0.88 [0.57, 1.36] .578 1.27 [0.76, 2.13] .359
Table 4. Multilevel regression models predicting the objective comprehensibility of explanations (the definition recognition task for RQ1).
Subjective Understanding of Explanations (RQ2)
In both studies, the reported level of subjective understanding was higher for the OT (Study 1: Mean (M) = 6.57, Standard Deviation (SD) = 2.37; Study 2: M = 6.73, SD = 2.21) than for the SC explanations (Study 1: M = 5.67, SD = 2.51; Study 2: M = 6.18, SD = 2.62). In comparison, the level of reported understanding of the OC explanations was comparable to that of the SC explanations in Study 1 (OC Study 1: M = 5.62, SD = 2.27), and higher than the reported understanding of SC explanations in Study 2 (OC Study 2: M = 6.81, SD = 2.13). The analyses revealed that when SC explanations were compared to all occlusion explanations, participants perceived the latter to be significantly more comprehensible than the former in both studies. However, the comparison of the different formats of the occlusion explanations to the SC explanations in separate analyses only found evidence for the superiority of OT over SC explanations. Moreover, in Study 1, participants found OT more comprehensible than OC explanations. The other comparisons were non-significant. Table 5 presents the results of the multilevel regression analyses.
Regression Model Study 1 Study 2
b p b p
 SC vs. OC & OT  0.46 [0.18, 0.74]  .001 0.58 [0.18 - 0.98]  .004
SC vs. OC 0.10 [-0.30, 0.49]  .637 0.51 [-0.07, 1.09]  .088
SC vs. OT 0.79 [0.41, 1.17]  ¡ .001  0.67 [0.18, 1.16]  .008
OC vs. OT 0.95 [0.10, 1.79]  .030 -0.09 [-0.97, 0.79]  .841
Table 5. Multilevel regression models predicting the difference in subjective understanding of explanations (RQ2).
Trustworthiness of Explanations (RQ3)
The reported level of trust was consistently the lowest for the SC explanations across the studies (Study 1: M = 4.95, SD = 2.45; Study 2: M = 6.73, SD = 2.21). The OT explanations scored the highest in Study 1 (Study 1: M = 5.47, SD = 2.32; Study 2: M = 5.66, SD = 2.18), and the OC explanations scored the highest in Study 2 (Study 1: M = 5.20, SD = 2.27; Study 2: M = 6.03, SD = 1.97). Similarly to the analyses of the subjective understanding of the explanations, when the occlusion explanations were compared with the SC explanations, participants trusted the occlusion explanations significantly more in both studies. However, the follow-up analyses revealed some contradictions between the two studies. Namely, Study 1 showed evidence for the superiority of OT over SC explanations, while Study 2 showed evidence for the preference for OC over SC explanations in terms of their trustworthiness. The other comparisons remained non-significant. Table 6 presents the results of the multilevel regression analyses.
Regression Model Study 1 Study 2
b p b p
SC vs. OC & OT 0.40 [0.13, 0.66]  .003 0.38 [0.02, 0.75]  .042
SC vs. OC 0.33 [-0.07, 0.73]  .107 0.57 [0.04, 1.10]  .035
SC vs. OT 0.44 [0.10, 0.78]  .012 0.23 [-0.25, 0.71]  .345
OC vs. OT 0.27 [-0.56, .1.10]  .530 0.37 [-0.48, 1.21]  .397
Table 6. Multilevel regression models predicting the difference in the trustworthiness of explanations (RQ3).
5. Discussion and Limitations
Across two studies with different samples (general population and medical students), we found a reliable preference for occlusion over Shapley value explanations. Our participants reported that they understood and trusted the occlusion explanations more than the SHAP explanations, as we expected. However, the comparison of all the explanations revealed that this effect may be solely driven by the fact that our participants found the OT explanations consistently more comprehensible than the SC explanations. Moreover, in Study 1, the OT explanations were also preferred to the SC explanations, though in Study 2 the results were less clear as there was no evidence that the former explanations were preferred over the latter. These findings are consistent with a notion of people preferring explanations in text over chart format (Bertrand et al., 2023), for instance, due to finding some graphs difficult to interpret (Durand et al., 2020; Galesic and Garcia-Retamero, 2011; Szymanski et al., 2021). This idea somewhat undermines the potential message that, in terms of content, FA explanations based on simple rather than complex models should be prioritised. Interestingly, the results regarding trustworthiness were slightly in conflict with this idea, since in our medical student sample we only found evidence for higher trust in the OC compared to the SC explanations. While these findings show some support for the role of explanation content, more research is needed to explore how people relate to and utilise different XAI techniques, and how and when an explanation’s format (e.g. text vs. chart) trumps the impact of its content. Future research should also explore the potential of tailoring explanation content and format to individual needs (Bhatt et al., 2020), such as their expertise in a field (Wang and Yin, 2023) or AI-systems (Schoeffer et al., 2022) and their skills to interpret numerical information (e.g. graph literacy (Galesic and Garcia-Retamero, 2011; Durand et al., 2020) or numeracy (Reyna et al., 2009; Gleaves et al., 2020)).
Our novel performance measure (i.e. the definition recognition task) could also not settle the question regarding the impact of content as we found no evidence that the participants were better at recognising the definition of either explanation. Nonetheless, the lack of positive findings could be related to how we measured performance. That is, the task of identifying the correct definition of explanations may have been too difficult as performance was only reliably over chance level for the occlusion explanations in the medical student sample potentially creating a floor effect for our measure. Future work should operationalise performance in other formats also.
Our studies carry some limitations. Both studies lacked a SHAP explanation in text format and so a complete design, making it difficult to fully separate the impact of the content of an explanation from the format. Nonetheless, as argued previously, it is not obvious how an ST explanations that is fairly represenative of SHAP can be created. Moreover, the relatively lower sample size of Study 2 may have caused some of our tests to be insensitive and non-significant preventing us from settling the question of whether occlusion explanations are perceived to be more comprehensible than SHAP explanations regardless of the format. For example, a significant difference between OC and SC explanations in terms of their comprehensibility would challenge the idea that preference for occlusion explanations is driven by a preference for charts. Finally, self-report measures are prone to biases, and a recent study (Nagendran et al., 2023) found no correlation between self-reported understanding of XAI and the impact of XAI on behaviour, casting some doubts on their generalisability to real-life behaviour. Future research should investigate this presumed dissociation and focus on behavioural measures of advice uptake (Pálfi et al., 2022; Pálfi et al., 2023; Panigutti et al., 2022) and understanding of advice.
Supplementary Material
Following are three examples of other vignettes used in the experiments.
Example 0.
Patient name: Rish Barrett (male), Age: 60, Current BMI: 37.55 (Height 175cm, Weight 115kg), Smoking: Smoker 5 cigarettes/day, Past medical history: Type 2 Diabetes Mellitus, Family history: Nil relevant.
Mr Barrett comes to see you for his results. He saw another GP at the surgery two weeks ago about the fact that after a long time he had lost a lot of weight within a short period of time (2 months) and could not understand why. Your colleague had organised some investigations. During your consultation, he reports that he has ongoing weight loss and has also been experiencing some abdominal pain over the past month. There are no other symptoms and examination findings are normal.
The cancer prediction algorithm estimates the risk of gastro-oesophageal cancer to be 2.53%.
Example 0.
Patient name: Natalie Clark (female), Age: 74, Current BMI: 34.60 (Height 170cm, Weight 100kg), Smoking: Ex-smoker, Past medical history: Nil, Family history: Nil relevant.
Natalie sees you today accompanied by her husband. Over the past 6 weeks, she has been experiencing heartburn and has been off her food. She has vomited on a few occasions and seen a bit of blood in the vomit. She also finds it difficult to swallow when eating. She has no other symptoms and examination findings are normal.
The cancer prediction algorithm estimates the risk of gastro-oesophageal cancer to be 40.24%.
Example 0.
Patient name: Marianne Foster (female), Age: 60, Current BMI: 27.85 (Height 163cm, Weight 74kg), Smoking: Never smoked, Past medical history: Nil, Family history: Nil relevant.
Mary comes back to see you for a follow-up consultation. Another GP at the surgery had organised some blood tests and a FIT test 4 weeks ago. On the electronic health record, you can see that during the last consultation she had complained of having lost her appetite as well as some weight (about 3kg over the previous 6 weeks) despite no changes in her diet or lifestyle. She reports that she is still off her food and has lost more weight. She has also noticed that it is becoming increasingly difficult for her to swallow when she tries to eat. The blood tests and FIT test which were organised came back normal. She denies any post-menopausal bleeding, has no other symptoms and examination findings are normal.
The cancer prediction algorithm estimates the risk of gastro-oesophageal cancer to be 8.27%.
Table 7 summarises the values they result in for the QCancer algorithm.
In the following, we give an example of the survey which was presented to users. First, users were asked for their consent:
We then asked the users the pre-test questions:
We then showed the users a demonstration of the main test questions:
For each set of vignette, prediction and explanation in the main test (which were as seen in the demonstration), we asked the following questions, with the last being on a new page.
Finally, we asked users the post-test questions: